OpenStudy (anonymous):

Solve the following problems using the sequence: five, thirteen, twenty-one, twenty-nine, thirty-seven, and so on a. Use words to write a rule for the sequence. b. Use the rule you wrote in the previous part to find the next three terms of the sequence. Be sure to show all your steps.

OpenStudy (jack1):

5, 13, 21, 29, 37... difference between each term = 8, 8, 8, 8... i think i see the pattern ;)

OpenStudy (anonymous):

ok

OpenStudy (jack1):

so... can you do the rest from here?

OpenStudy (anonymous):

so what would i say

OpenStudy (jack1):

u tell me... they want u to use words (not numbers) to explain the pattern... including the starting term

OpenStudy (triciaal):

there are 2 types of sequences: arithmetic with a common difference and geometric with a common ratio

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

ok

OpenStudy (jack1):

so how do u think u should explain it?

OpenStudy (anonymous):

i have noo clue

OpenStudy (jack1):

k... what's the first term?

OpenStudy (triciaal):

read again what jack1 wrote originally

OpenStudy (anonymous):

5 ?

OpenStudy (anonymous):

8

OpenStudy (jack1):

yep, perfect so something like "the sequence starts at 5..." then... what's the pattern/sequence?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

it goes up by 8

OpenStudy (jack1):

perfect, so "the sequence begins at 5... then increases by 8... with each new term"

OpenStudy (anonymous):

ok thank you sooo much

OpenStudy (jack1):

all good now part b what are the next 3 terms in the sequence (after 37?)

OpenStudy (anonymous):

ok

OpenStudy (usukidoll):

37+ 8 =?

OpenStudy (jack1):

no... srsly... what are the next 3 terms? @lauren.ascough

OpenStudy (anonymous):

um

OpenStudy (anonymous):

37 + 8 = 46 @UsukiDoll

OpenStudy (usukidoll):

close there's a typo

OpenStudy (anonymous):

what do you mean

OpenStudy (usukidoll):

37+8 is actually 45

OpenStudy (anonymous):

ok sorry

OpenStudy (usukidoll):

it's ok :) so we got one term... two more terms to go and since we are increasing by 8 all the time we solve 45+8 = ?

OpenStudy (anonymous):

45+8=53

OpenStudy (usukidoll):

yup... one more term 53+8 = ?

OpenStudy (jack1):

5, 13, 21, 29, 37... 1st term = 5 2nd term = 5 + 8 = 13 3rd term = 5 + 8 + 8 = 21 4th term = 5 + 8 + 8 + 8 = 29 5th term = 5 + 8 + 8 + 8 + 8 = 37 6th term = ??? 7th term = ??? 8th term = ???

OpenStudy (anonymous):

53+8=61

OpenStudy (usukidoll):

yes... so our 3 terms are 45, 53, 61

OpenStudy (anonymous):

ok

OpenStudy (usukidoll):

6th term = 5+8+8+8+8+8 = 45 7th term = 5+8+8+8+8+8+ 8 = 53 8th term = 5+8+8+8+8+8+ 8 + 8 =61

OpenStudy (usukidoll):

so that's about it for this part.

OpenStudy (anonymous):

ok so what do i write for my final answer

OpenStudy (usukidoll):

Our sequence is increasing by 8 and by following that pattern, we retrieve the next three terms which is 45,53, and 61

OpenStudy (anonymous):

ok thank you for your help